OpenStudy (anonymous):

What are the zeros of the function? y = (x+3) (x-3) (x+4)

OpenStudy (anonymous):

solve each perenthetical group one by one set it to zero

OpenStudy (anonymous):

so (x+3) =0 -3 (x-3) =0 3 (x+4) =0 -4